# EPIC-SOUNDS: A Large-Scale Dataset of Actions that Sound

### v2

@@ VI-E Audio-Visual Sound Detection
Training & Validation Configuration. The visual models are trained in the same way as their audio-counterparts, with visual input features. To create a single set of proposals, we combine the predictions of each time step from the audio and visual models. We follow [34] and re-weight the confidence of the proposal and its action boundaries of each proposal by:
@@
Results. We also report audio-visual detection results in Table IV. In contrast to audio-only, TriDet-AV now outperforms ActionFormer-AV on the validation set by 0.3 average mAP, whereas ActionFormer-AV outperforms by 0.6 and 0.4 average mAP on the detection test and entire test set respectively. In comparison to their audio-only version, TriDet-AV sees an additional boost of 1.5, 0.4 and 0.5 average mAP on the validation, detection and entire test set and ActionFormer-AV exhibits an increase of 1.0, 1.1 and 1.0 average mAP for the same splits. The additional performance boost of both models when integrating the visual modality highlights the beneficial information shared between them. Again, we show qualitative results for the audio-visual extension of the baselines in Figure 12. We see that the inclusion of the visual modality can help eliminate false positive predictions as well as improve the regressed boundaries.
@@ VII Summary and Impact
Impact. Following the introduction of EPIC-SOUNDS in 2023, a number of works have built on this dataset. We summarise these works here. [38] introduces a method for automatically producing audio-centric captions from video-text datasets using an LLM. They use the audio labels provided by EPIC-SOUNDS as prompts to generate audio descriptions. [39] proposes a self-supervised approach to learn how actions sound. To achieve this, they use a novel embedding to reinforce correlations between audio, video, and text modalities. Their method is evaluated on the EPIC-SOUNDS recognition challenge. [40] utilises a method inspired by latent diffusion models for video-to-audio synthesis, to generate high-quality, synchronised audio. They fine-tune and evaluate their method on EPIC-SOUNDS  showcasing its ability to generate accurate audios for samples such as ‘open drawer’ and ‘plate clinking’. [41] propose a multi-modal, autoregressive model jointly modelling audio-visual information and balancing the learning between the two modalities to produce efficient representations. They evaluate their model on the EPIC-SOUNDS validation set, producing current SOTA recognition results (79.4%). The previous arXiv version is available here : https://arxiv.org/abs/2302.00646v1.

### v3

@@ VI-E Audio-Visual Sound Detection
Training & Validation Configuration. The visual models are trained in the same way as their audio-counterparts, with visual input features. To create a single set of proposals, we combine the predictions of each time step from the audio and visual models.
We follow [35] and re-weight the confidence $\mathbf{p}(\cdot)$ and action boundaries $\mathbf{d}(\cdot)$ of each proposal by:
@@
Results. We also report audio-visual detection results in Table V. In contrast to audio-only, TriDet-AV now outperforms ActionFormer-AV on the validation set by 0.3 average mAP, whereas ActionFormer-AV outperforms by 0.6 and 0.4 average mAP on the detection test and entire test set respectively. In comparison to their audio-only version, TriDet-AV sees an additional boost of 1.5, 0.4 and 0.5 average mAP on the validation, detection and entire test set and ActionFormer-AV exhibits an increase of 1.0, 1.1 and 1.0 average mAP for the same splits. The additional performance boost of both models when integrating the visual modality highlights the beneficial information shared between them. Again, we show qualitative results for the audio-visual extension of the baselines in Figure 13. We see that the inclusion of the visual modality can help eliminate false positive predictions as well as improve the regressed boundaries.
@@ VII Summary and Impact
Impact. Following the introduction of EPIC-SOUNDS in 2023, a number of works have built on this dataset. We summarise these works here. [39] introduces a method for automatically producing audio-centric captions from video-text datasets using an LLM. They use the audio labels provided by EPIC-SOUNDS as prompts to generate audio descriptions. [40] proposes a self-supervised approach to learn how actions sound. To achieve this, they use a novel embedding to reinforce correlations between audio, video, and text modalities. Their method is evaluated on the EPIC-SOUNDS recognition challenge. [41] utilises a method inspired by latent diffusion models for video-to-audio synthesis, to generate high-quality, synchronised audio. They fine-tune and evaluate their method on EPIC-SOUNDS  showcasing its ability to generate accurate audios for samples such as ‘open drawer’ and ‘plate clinking’. [42] propose a multi-modal, autoregressive model jointly modelling audio-visual information and balancing the learning between the two modalities to produce efficient representations. They evaluate their model on the EPIC-SOUNDS validation set, producing current SOTA recognition results (79.4%).